OpenStudy (anonymous):

Write the complex number in the form a + bi. (5 points) 6(cos 330° + i sin 330°)

OpenStudy (anonymous):

If somebody could just walk me through it please and thank you

OpenStudy (primeralph):

Find cos330 and sin300, then expand.

OpenStudy (anonymous):

6 ( (.8660254038) + i (-.5))

OpenStudy (anonymous):

5.196152423 - 3i

OpenStudy (zzr0ck3r):

use the exact value, you should be able to find this at this point \[cos(330) = \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

so then it's\[6(\sqrt{3}/2 +-.5i)\]

OpenStudy (primeralph):

Yes. Good job.

OpenStudy (primeralph):

Expand it though.

OpenStudy (anonymous):

\[3 \sqrt{3} - 3i\]

OpenStudy (primeralph):

Yeah.

OpenStudy (anonymous):

yeah, thank you all!